OpenStudy (australopithecus):

Integration by parts problem, \[\int\limits_{}^{}e^{2x}\sin(3x)dx\] Trying to relearn integration, very rusty on it, haven't taken a calc course in almost 2 years. Totally stuck with this problem, any hints or help would be appreciated

OpenStudy (compphysgeek):

You remember how integration by parts works? \[\int u' v dx = uv - \int u v' dx\] Here you probably will have to apply the rules twice so that you'll have something like \[ I = uv - (rs + I) \] then you solve for \(I\) to get the solution.

OpenStudy (australopithecus):

I remember the rule just fine, but aren't these just functions that repeat endlessly? \[\int\limits_{}^{} \sin(3x)e^{2x}dx = \frac{\sin(3x)e^{2x}}{2} + \int\limits \frac{e^{2x}\cos(3x)}{6}\]

OpenStudy (compphysgeek):

No, because you'll get an integral of the form \[- k \int e^{2 x} \sin(3 x} dx \] if you apply integration by parts again. It's the same integral with which you started so that you can solve for that integral as I tried to describe in my first answer.

zepdrix (zepdrix):

Woops, you differentiated the sin(3x) so you get a 3 on top, not on bottom:\[\int\limits\limits_{}^{} \sin(3x)e^{2x}dx = \frac{\sin(3x)e^{2x}}{2} + \int\limits\limits \frac{\color{red}{3}e^{2x}\cos(3x)}{2}\]Integrate one more time. Make sure you assign your exponential to be your dv again, otherwise you'll end up undoing your work from before.

OpenStudy (agentjamesbond007):

When memorising the formula for UV-substitution (Integration by Parts), use the mneumonic "UltraViolet-VooDu" UV-∫vdu

zepdrix (zepdrix):

lol clever :3

OpenStudy (australopithecus):

I already have it memorized but that is pretty clever

OpenStudy (australopithecus):

I noticed my mistake Im just tired

OpenStudy (australopithecus):

As for settng exponential as du do you mean setting u = e^2x that will just give me du = 2e^2xdx

OpenStudy (compphysgeek):

in the first step you set \(e^{2x}\) as your \(v'\). you need to make sure to do that again

zepdrix (zepdrix):

no no. :O The first time you integrated it looks like you assigned: u=sin(3x), dv=e^(2x) So it's important that you do likewise again: u=cos(3x), dv=e^(2x)

OpenStudy (australopithecus):

\[\frac{\sin(3x)e^{2x}}{2} + \frac{e^{2x}3\cos(3x)}{4} + \int\limits\limits 9\sin(3x)\frac{e^{2x}}{4}dx\]

OpenStudy (australopithecus):

oh I think I get what you are saying, set u = cos(3x) du = -3sin(3x)dx then I just go arccos(u)/3 = x then I just sub that into e^2x

OpenStudy (australopithecus):

wait no that will make things even more complicated blah

zepdrix (zepdrix):

\[\int\limits\limits\limits_{}^{} \sin(3x)e^{2x}dx = \frac{\sin(3x)e^{2x}}{2} \color{red}{-} \int\limits\limits\limits \frac{\color{red}{3}e^{2x}\cos(3x)}{2}\]Oh I just noticed, shouldn't this first sign have been minus? You differentiated sin(3x) to positive 3cos(3x), yes?

OpenStudy (australopithecus):

yeah you are right I made a mistake there

zepdrix (zepdrix):

\[\large\rm \color{orangered}{\int\limits\sin(3x)e^{2x}dx}=\frac{1}{2}\sin(3x)e^{2x}-\frac{3}{4}\cos(3x)e^{2x}+\frac{9}{4}\color{orangered}{\int\limits\sin(3x)e^{2x}dx}\]Ok so by integrating twice you've gotten to this point.

zepdrix (zepdrix):

If you call it something else, it might be easier to work with. \[\large\rm \color{orangered}{\mathcal I}=\frac{1}{2}\sin(3x)e^{2x}-\frac{3}{4}\cos(3x)e^{2x}+\frac{9}{4}\color{orangered}{\mathcal I}\] It's just algebra from this point. Solve for I.

OpenStudy (australopithecus):

I- 9/4I = 5/4I so (4/5)*( 1/sin(3x)e^(2x) - 3/4cos(3x)e^2x )

zepdrix (zepdrix):

Yes, you've got the right idea. Sec, I wanna check your signs a minute...

OpenStudy (australopithecus):

man Im so tired my signs are going to always be messed up lol, I totally dont remember this trick at all.

OpenStudy (australopithecus):

pretty neat though

zepdrix (zepdrix):

Oh oh, remember we fixed that minus sign a little later on? So we need to distribute that negative to each term in the `by parts` application. So the 9/4I term should be negative.

zepdrix (zepdrix):

in the second `by parts` application*

zepdrix (zepdrix):

\[\large\rm \color{orangered}{\mathcal I}=\frac{1}{2}\sin(3x)e^{2x}-\frac{3}{4}\cos(3x)e^{2x}-\frac{9}{4}\color{orangered}{\mathcal I}\]So easy to lose track of these things :( lol

OpenStudy (australopithecus):

oh oh damns well its the same principle I'm way too tired at the moment to finish this thanks for showing me this trick though, I will just do it again tomorrow and check it on wolfram thanks for the help

zepdrix (zepdrix):

heh :3